OpenStudy (anonymous):

A program to simulate coin tosses is used to simultaneously toss 1,000 coins. If the number of heads displayed in 5 trials of the simulation total 528, 486, 492, 545, and 505, what is the experimental probability of getting heads on a coin toss? Round your answer to the nearest percent. 50% 51% 49% 53%

OpenStudy (anonymous):

Is there anyone that's good a prob. and stats that can help me?

OpenStudy (anonymous):



OpenStudy (anonymous):

Thanks, I already answered the question but I didn't get it correct. Do you think you can help me with 5 problems, John?

OpenStudy (anonymous):

I can attempt, yes.

OpenStudy (anonymous):

Omg thank you! Hold on a second and I will type them :D

OpenStudy (anonymous):

A die of unknown bias is rolled 20 times, and the number 3 comes up 6 times. In the next three rounds (the die is rolled 20 times in each round), the number 3 comes up 6 times, then 5 times, and finally 7 times. Which experimental probability is consistent with this simulation?

OpenStudy (anonymous):

The probability of 3 is 15% The probability of 3 is 30% The probability of 3 is 45% The probability of 3 is 60%

OpenStudy (anonymous):

You will probably need the average of the numbers it was rolled. So \[\frac{ (6+6+5+7) }{ 4 } = \frac{ 24 }{ 4 } = 6\]

OpenStudy (anonymous):

6 is what percent of 20?

OpenStudy (anonymous):

30?

OpenStudy (anonymous):

\[6 = \frac{ x }{ 100 } \times 20\]

OpenStudy (anonymous):

Correct so the probability of attaining a 3 should be 30%

OpenStudy (anonymous):

Thank you! Okay! Next question.. A person playing a game of chance has a 0.25 probability of winning. If the person plays the game 20 times and wins half of that number of times, what is the difference between the theoretical probability and the experimental probability of that person winning ?

OpenStudy (anonymous):

0.50 0.25 0.30 0.75

OpenStudy (anonymous):

You know the probability of winning is 0.25 When the person played 20 times they won half the time so,\[\frac{ 10 }{ 20 } = \frac{ 1 }{ 2 }= 0.50\] Now we are asked to find the "difference" between the theoretical(0.25) and the experimental (0.50).

OpenStudy (anonymous):

Difference refers to subtraction of one term by another and in this case we know both terms.

OpenStudy (anonymous):

So it's .50?

OpenStudy (anonymous):

Nope. We need to subtract one term from the other.

OpenStudy (anonymous):

Okay so we subtract .50-.25? that equals .25

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

Thanks! Okay next question :D A simulated game of chess is programmed between two computers. The game is supposed to be biased in favor of player B winning 4 out of 5 times. Which is the most suspicious set of outcomes of 30 games played between the two computers?

OpenStudy (anonymous):

AABABBABBBBBBBBABBBABBBBBBBBA ABBABBAABBBBBABBBBBBBBBBBBBABB ABBABABABABBABABABABABABABAA ABBABBBABBBBBBBBABBBA

OpenStudy (anonymous):

B should win 4/5 times or,\[\frac{ 4 }{ 5 } = 0.80\]

OpenStudy (anonymous):

a. AABABBBABBBBBBBBABBBABBBBBBBBA b. ABBABBAABBBBBABBBBBBBBBBBBBABB c. ABBABABABABBABABABABABABABABAA d. ABBABBBABBBBBBBBABBBABBBBBBBBB ^sorry, these are the correct choices

OpenStudy (anonymous):

If you add up the total number of times B won and divide it by the 30 games played. It should be 4/5 or 0.80.

OpenStudy (anonymous):

So it's B? Are you sure?

OpenStudy (anonymous):

I am not sure, i have not performed the operation.

OpenStudy (anonymous):

I looked the question up and the person said it's not b. It's something else.

OpenStudy (anonymous):

It is not B, B wins 24 times out of 30 or, \[\frac{ 24 }{ 30 } = \frac{ 12 }{ 15 } = \frac{ 4 }{ 5 } = 0.80\]

OpenStudy (anonymous):

Did not say it was B, I was referring to "player B" before.

OpenStudy (anonymous):

Oh, I thought you were saying it was B XD. Okay, so B won four out of 5 times...and you divide it by 30? O.o

OpenStudy (anonymous):

No B won 24/30 times and i just simplified, hence the = signs above.

OpenStudy (anonymous):

24/30 = 0.80 in your calculator

OpenStudy (anonymous):

Okay, I got .80 on my calculator.

OpenStudy (anonymous):

So it is correct and does not look suspicious.

OpenStudy (anonymous):

You are looking for the "most" suspicious which means the one that is way different from your 4/5 or 0.80

OpenStudy (anonymous):

I don't think it looks suspicious.

OpenStudy (anonymous):

It doesn't at all. It is exactly 0.80 so it is not our suspicious game.

OpenStudy (anonymous):

Okay. So how do we find the one that's the most suspicious?

OpenStudy (anonymous):

Find the number of times B won in each game and divide it by the total number of games played.

OpenStudy (anonymous):

a. AABABBBABBBBBBBBABBBABBBBBBBBA, A= 7 , B = 23 b. ABBABBAABBBBBABBBBBBBBBBBBBABB, A = 6, B = 24 c. ABBABABABABBABABABABABABABABAA, A= 15, B =15 d. ABBABBBABBBBBBBBABBBABBBBBBBBB, A = 5 , B = 25

OpenStudy (anonymous):

Now which looks most suspicious?

OpenStudy (anonymous):

Okay, so A) B won 22 times? C)14 and D) 26 times?

OpenStudy (anonymous):

I got different values than you did?

OpenStudy (anonymous):

We both are close though even with both our values which of them looks suspicious? And where B does not win 4/5 times

OpenStudy (anonymous):

B meaning Player B

OpenStudy (anonymous):

C looks suspicious to me. I probably didn't count the B's correctly.

OpenStudy (anonymous):

It is alright. Correct because in my count B won half the time not close to 4/5. And in your count B won almost half the time which is also not close to 4/5.

OpenStudy (anonymous):

So it is C? :D

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

C is the most suspicious game with 0.50 probability of winning which is farther from 0.80 compared to the other games.

OpenStudy (anonymous):

Thank you! :D okay, just 2 more questions! A sharpshooter in a training simulation has a miss rate of 1%. If H represents a hit and M represents a miss, which is the most likely valid result of a simulation model of 30 practice shots?

OpenStudy (anonymous):

.HHHHHHHHHHHHHHHHHHHHHHHHHHHHHH b. MHHHHHMHHHHHHHHHHHHHMHHHHHHHHH c.HHHHHHHMHHHHHMHHHHHHMHHHHHHMHH d.HMHHHHHHHHHHHHHHHHHHHHMHHHHHHH

OpenStudy (anonymous):

*a

OpenStudy (anonymous):

Most likely "valid" result so we are looking for the results that most depicts a 1% miss rate.

OpenStudy (anonymous):

Looks like A to me....

OpenStudy (anonymous):

First we should ask: What is 1% of 30 games?

OpenStudy (anonymous):

.30?

OpenStudy (anonymous):

Correct it is less than 1 so we are looking for when he is least likely to miss.

OpenStudy (anonymous):

Okay :D so do we have to count the H's or M's?

OpenStudy (anonymous):

You are correct all of other games have a higher miss rate than the game A which has a miss rate of 0% closest to 1%

OpenStudy (anonymous):

I'm pretty sure it's a....

OpenStudy (anonymous):

That is correct.

OpenStudy (anonymous):

You are correct all of other games have a higher miss rate than the game A which has a miss rate of 0% closest to 1%

OpenStudy (anonymous):

Thanks! One more :D A soccer team wins 65% of its matches, and 15% of its matches end in a draw. If the team is scheduled to play 20 matches, about how many matches is it expected to lose?

OpenStudy (anonymous):

Well you know that A soccer team will win 65% and 15% will be draws so out of 100% the rest must be loses.